Air is contained in a vertical piston-cylinder assembly fitted with an electrical resistor. The atmospheric pressure is 100 kPa and point has a mass of 50 kg and a face area of 0.1 m2. Electric current passes through the resistor, and the volume of of air slowly increases by 0.045 m3. The mass of the air is 0.3 kg, and its specific energy increases by 42.2 kJ/kg. Assume the assembly ( including the piston) is insulated and neglect the friction between the cylinder and piston, g = 9.8 m/s2. Determine the heat transfer from the resistor to air for a system consisting of (a) the balloon, (b) the air and the piston


1
Expert's answer
2021-04-13T06:38:08-0400



Assumptions:

1. Two closed systems are under consideration, as shown in schematic.

2. The only heat transfer is from the resistor to the air. AKE = APE=0 (for air)

3. The internal energy of the piston is not affected by the heat transfer.




a) Taking the air as the system,

(∆KE + ∆PE + ∆U) = Q - W

Q=W+ ∆Uair

For this system work is done at the bottom of the piston. The work done by the system is:

"W = \int^{V_2}_{V_1}pdV = p (V_2-V_1)" ------- (at constant pressure)


The pressure acting on the air can be found from:


"PA_{piston }= m_{pistong}g+ P_{atm}A_{piston}"


"P =\dfrac{m_{piston}g}{A_{piston}}+ P_{atm}"


"P = \dfrac{50× 9.81}{0.1} + 100kPa"


P = 104.91 kPa


Thus, the work is

W = (104.91 kPa)(0.045m³) = 4.721kJ


With ∆Uair = mair ∆uair the heat transfer is;

Q =W + mair ∆uair = 4.721 kJ + (0.3 kg)(42.2 kJ/kg) = 17.38 kJ




b) system consisting the air and the piston. The first law becomes:


(∆KE + ∆PE + ∆U)air + (∆KE + ∆PE + ∆U)piston = Q - W,

where (∆KE = ∆PE)air = 0 and (∆KE = ∆U)piston = 0.

Thus, it simplifies to;

(∆U)air + (∆PE)piston = Q - W


For this system, work is done at the top of the piston and pressure is the atmospheric pressure. The work becomes

W = Patm∆V = (100 kPa)(0.045m') = 4.5 kJ


The elevation change required to evaluate the potential energy change of the piston can be found from the volume change:

∆z = ∆V/Apiston = 0.045 m³/0.1 m² = 0.45 m


(∆PE =mpistong∆z = (50 kg)(9.81 m/s')(0.45 m) = 220.73 J = 0.221 kJ


Q = W + (∆PE)piston + mair∆uair

Q=4.5 kJ + 0.221 kJ + (0.3 kg)(42.2 kJ/kg) = 17.38 kJ



Note that the heat transfer is identical in both systems.